OpenStudy (kewlgeek555):

The art museum had a total of 224 visitors on Tuesday. Visitors older than 18 paid $12 for admission. Visitors 18 years or younger paid $8 for admission. The museum collected a total of $2,520. How many visitors were 18 or younger? @agent0smith @mathstudent55 @ganeshie8 @thomaster @terenzreignz @hero

OpenStudy (kewlgeek555):

So this is what I did: I created a system of equations: {8u + 12o = 2520 {u + o = 224

OpenStudy (anonymous):

well I divided and I think its 315

OpenStudy (kewlgeek555):

Now, I am not sure if I just use the Elimination Method or Substituting. I got all confused with the elimination method, so I just substituted and this is what I did:

OpenStudy (kewlgeek555):

8u + 12o = 2520 u + o = 224 - o = u - 224 8u + 12(u - 224) = 2520 8u + 12u - 2688 = 2520 20u - 2688 = 2520 20u = 5208 u = 260.4

OpenStudy (kewlgeek555):

I forgot to say that u means 18 & under and o means over 18.

OpenStudy (kewlgeek555):

AFK

OpenStudy (anonymous):

42 people were under 18, do you want an explanation?